Calculate K for the following reaction:

A + 2 B ⇌ 2 C + D

given the following K values:

A ⇌ D + 2 E K1 = 72.41

C ⇌ B + E K2 = 38.81
1
Expert's answer
2020-02-17T07:09:02-0500

Solution.

K1 and K2 equally:


"K1=\dfrac{[D]*[E]^2}{[A]}" ;

"K2=\dfrac{[B]*[E]}{[C]}" .

Express K through K1 and K2:

"K= \\dfrac{[C]^2*[D]}{[A]*[B]^2}= \\dfrac {[C]^2*[D]*[E]^2}{[A]*[B]^2*[E]^2}=\\dfrac {[D]*[E]^2}{[A]}*\\dfrac {[C]^2} {[B]^2*[E]^2}=K1*\u200b\t\\Bigg ( { \\dfrac {1}{K2}}\\Bigg)^2=\u200b\t\\dfrac{K1}{K2^2}" .


Calculate K:


"K=\u200b\t\n \\dfrac{72.41}{38.81^2}=0.048."


Answer: 0.048.